Mickley Corporation produces two products, Alpha6s and Zeta7s, which pass through two operations, Sintering and Finishing. Each of the products uses two raw materials—X442 and Y661. The company uses a standard cost system, with the following standards for each product (on a per unit basis): Raw Material Standard Labor Time Product X442 Y661 Sintering Finishing Alpha6 1.8 kilos 2.0 liters 0.20 hours 0.80 hours Zeta7 3.0 kilos 4.5 liters 0.35 hours 0.90 hours Information relating to materials purchased and materials used in production during May follows: Material Purchases Purchase Cost Standard Price Used in Production X442 14,500 kilos $52,200 $3.50 per kilo 8,500 kilos Y661 15,500 liters $20,925 $1.40 per liter 13,000 liters The following additional information is available: The company recognizes price variances when materials are purchased. The standard labor rate is $19.80 per hour in Sintering and $19.20 per hour in Finishing. During May, 1,200 direct labor-hours were worked in Sintering at a total labor cost of $27,000, and 2,850 direct labor-hours were worked in Finishing at a total labor cost of $59,850. Production during May was 1,500 Alpha6s and 2,000 Zeta7s. Required: 1. Complete the standard cost card for each product, showing the standard cost of direct materials and direct labor. 2. Compute the materials price and quantity variances for each material. 3. Compute the labor rate and efficiency variances for each operation.
Answer #1

Part 1.

Standard
Quantity or Hours
Standard
Price or Rate
Standard Cost
Alpha6:
Direct materials-X442 1.8 kilos 3.50 per kilo 6.30
Direct materials-Y661 2.0 liters 1.40 per liter 2.80
Direct labor-Sintering 0.20 hours 19.80 per hour 3.96
Direct labor-Finishing 0.80 hours 19.20 per hour 15.36
Total $ 28.42
Zeta7:
Direct materials-X442 3.0 kilos 3.50 per kilo 10.50
Direct materials-Y661 4.5 liters 1.40 per liter 6.30
Direct labor-Sintering 0.35 hours 19.80 per hour 6.93
Direct labor-Finishing 0.90 hours 19.20 per hour 17.28
Total $ 41.01

Part 2.

The computations to follow will require the standard quantities allowed

for the actual output for each material.

Standard Quantity Allowed

Material X442:

Production of Alpha 6 (1.8 kilos per unit ×1,500 units)......2,700 kilos

Production of Zeta 7 (3.0 kilos per unit×2,000 units)........6,000 kilos

Total..............................................................................8,700 kilos

Material Y661:

Production of Alpha 6(2.0 liters per unit×1,500 units).....3,000 liters

Production of Zeta 7(4.5 liters per unit×2,000 units).......9,000 liters

Total..............................................................................12,000 liters

Direct Materials Variances

Material X442:

Materials quantity variance = SP (AQ–SQ)= $3.50 per kilo (8,500 kilos–8,700 kilos)= $700 F

Materials price variance = AQ (AP–SP)= 14,500 kilos ($3.60 per kilo*–$3.50 per kilo)= $1,450 U

*$52,200÷14,500 kilos = $3.60 per kilo

Direct Materials Variances—

Material Y661:

Materials quantity variance = SP (AQ–SQ)= $1.40 per liter (13,000liters–12,000 liters)= $1,400 U

Materials price variance = AQ (AP–SP)= 15,500 liters ($1.35per liter*–$1.40 per liter)= $775F

*$20,925÷15,500 liters = $1.35per liter

Part 3.

The computations to follow will require the standard quantities allowed for the actual output for direct labor in each department.

Standard Hours Allowed

Sintering:

Production of Alpha6(0.20 hours per unit×1,500 units).....300 hours

Production of Zeta7(0.35 hours per unit×2,000 units).......700 hours

Total..................................................................................1,000 hours

Finishing:

Production of Alpha6(0.80 hours per unit×1,500 units)....1,200 hours

Production of Zeta7(0.90 hours per unit×2,000 units)......1,800 hours

Total..................................................................................3,000 hours

Direct Labor Variances—Sintering:

Labor efficiency variance = SR (AH–SH)= $19.80 per hour (1,200 hours–1,000 hours)= $3,960 U

Labor rate variance = AH (AR–SR)= 1,200 hours ($22.50 per hour*–$19.80 per hour)= $3,240 U

*$27,000÷1,200 hours = $22.50 per hour

Direct Labor Variances—Finishing:

Labor efficiency variance = SR (AH–SH)= $19.20 per hour (2,850 hours–3,000 hours)= $2,880 F

Labor rate variance = AH (AR–SR)= 2,850 hours ($21.00 per hour*–$19.20 per hour)= $5,130 U

*$59,850÷2,850 hours = $21.00 per hour